REU: TENTATIVE GRADUATE STUDENT ASSIGNMENTS

The calendar of the program, with duties of undergraduates is:

Week 1 (June 23 – 27) Orientation and YSP training sessions Week 2 (June 30 – July 3) YSP duties 9 a.m. - 2:30 p.m. Week 3 (July 7 – 11) YSP duties 9 a.m. - 2:30 p.m. Week 4 (July 14 – 18) YSP duties 9 a.m. - 2:30 p.m. Week 5 (July 21 – 25) YSP duties 9 a.m. - 2:30 p.m. Week 6 (July 28 – August 1) SESAME duties either morning or afternoon Week 7 (August 4 – 8) SESAME duties either morning or afternoon Week 8 (August 11 – 15)

Tentative assignments: feel free to complain or make alternative requests. There are many factors that may lead to later changes: attendance at the various courses, developing mentor relationships, DRP, etc.

From Mathematics: *Mohammed Abouzaid [Farb–Hruska, May] *Jayadev Athreya [Hruska, Constantin] *David Balduzzi [Babai (weeks 1–4), Fefferman, Kirr, Santosa] *Angela Barnhill [Herrmann] *Andrew Blumberg [Fefferman, Santosa, May] *Jeff Clouse [Fefferman, Kirr, Santosa, May] *Jeremy Copeland [Babai, Constantin, Presentations] *Matthew Day [Farb, May] *Nick Gurski [Farb–Hruska, May, Lewicka] *Gautam Iyer [Constantin, Lewicka] *Sanjeevi Krishnan [fefferman, Kirr, Santosa, May] *Craig Jackson [Farb–Hruska, Lewicka] *Ben Lee [Herrman, Santosa, May] *Sharon McCathern [Herrmann] *Mridul Mehta [Kirr, Santosa, Babai (weeks 5–8)] *Courtney Morris [Farb–Hruska, May, Presentations] *Ann Scheels [Fefferman, Kirr, Santosa, Constantin] *Kacey Walker [Herrmann, Santosa, Constantin, Presentations] *Ben Wieland [Babai (weeks 5–8)] *Dani Zarnescu [Constantin, Lewicka]

From Computer Science: *Ivona Bezakova [Babai] *Varsha Dani [Babai] *Tom Hayes [Babai] *Daniel Stefankovic [Babai] List of courses with graduate student assignments:

1. DISCRETE MATHEMATICS, Laci Babai (weeks 1–8)

The first module (weeks 1–4) will focus on the interaction between linear algebra, combinatorics, and algorithms. The second module (weeks 5–8) will focus on combinatorial and algorithmic aspects of finite groups.

Ivona Bezakova, weeks 1–8 Varsha Dani, weeks 1–8 Tom Hayes, weeks 1–8 Daniel Stefankovic, weeks 1–8 David Balduzzi, weeks 1–4 Jeremy Copeland, weeks 1–4 Mridul Mehta, weeks 5–8 Ben Wieland, weeks 5–8

2 KNOTS and LINKS, Benson Farb (weeks 1–2), Chris Hruska (weeks 3–4) Mohammed Abouzaid Jayadev Athreya (weeks 3–4) Matthew Day (weeks 1–2) Nick Gurski Craig Jackson Courtney Morris

3. INTRODUCTION TO GROUPS AND GEOMETRY, Diane Herrmann (weeks 1–2) Angela Barnhill Ben Lee Sharon McCathern Kacey Walker

4. INVITATION TO PROBABILITY THEORY, Robert Fefferman (week 1) David Balduzzi Andrew Blumberg Jeff Clouse Sanjeevi Krishnan Ann Scheels

5. TOPICS IN ODE’S, Eduard Kirr (week 2) David Balduzzi Jeff Clouse Sanjeevi Krishnan Mridul Mehta Ann Scheels

6. MATHEMATICS IN INDUSTRIAL APPLICATIONS, Fadil Santosa (weeks 3–4)

David Balduzzi Andrew Blumberg Jeff Clouse Ben Lee Mridul Mehta Sanjeevi Krishnan Ann Scheels Kacey Walker

7. FOLLOW UP ON AN INTIVATION TO PROBABILITY THEORY: MARKOV CHAINS, MARTINGALES, AND MORE …, Peter Constantin (weeks 5–6) Jayadev Athreya Jeremy Copeland Gautam Iyer Ann Scheels Kacey Walker Dani Zarnescu

8. FINITE TOPOLOGICAL SPACES, Peter May (weeks 5–7) Mohammed Abouzaid Andrew Blumberg Jeff Clouse Matthew Day Nick Gurski Sanjeevi Krishnan Ben Lee Courtney Morris

9. INTRODUCTION TO TOPOLOGICAL DEGREE IN EUCLIDEAN SPACES, Marta Lewicka (weeks 7–8) Nick Gurski Gautam Iyer Craig Jackson Dani Zarnescu

10. STUDENT PRESENTATIONS (week 8) Organizers: Jeremy Copeland Courtney Morris Kacey Walker